### v3

@@ 4.3 Distributional misspecification
A model is distributionally misspecified when all observations in a given sample are generated by a distribution that is different from the model distribution. In the context of the polychoric model, this means that all ordinal observations are generated by a latent distribution that is nonnormal. Let $G$ denote the unknown nonnormal distribution that the latent variables $(\xi,\eta)$ jointly follow under distributional misspecification. The object of interest is the population correlation between latent $\xi$ and $\eta$ under distribution $G$, for which the normality-based ML estimator of Olsson, (1979) turns out to be substantially biased in many cases (e.g., Lyhagen & Ornstein,, 2023; Foldnes & Grønneberg,, 2020, 2022; Jin & Yang-Wallentin,, 2017). As such, distributional misspecification is fundamentally different from partial misspecification: In the former, one attempts to estimate the population correlation of the nonnormal and unknown distribution that generated a sample, instead of estimating the polychoric correlation coefficient (which is the correlation under standard bivariate normality). In the latter, one attempts to estimate the polychoric correlation coefficient with a sample that has only been partly generated by latent normality (that is, the polychoric model). The assumption that the polychoric model is only partially misspecified for some uninformative observations enables one to still estimate the polychoric correlation coefficient of that model, which would not be feasible under distributional misspecification (at least not without additional assumptions).
Despite not being designed for distributional misspecification, the robust estimator introduced in the next section (Section 5) can offer a robustness gain in some situations where the polychoric model is distributionally misspecified. We discuss this in more detail in Section 8.
@@ 5 Robust estimation of polychoric correlation
The consistency of ML estimation of an arbitrary model crucially depends on correct specification of that model. Indeed, ML estimation can be severely biased even when the assumed model is only slightly misspecified (e.g., Huber & Ronchetti,, 2009; Hampel et al.,, 1986; Huber,, 1964). For instance, in many models of continuous variables like regression models, one single observation from a different distribution can be enough to make the ML estimator converge to an arbitrary value (Huber & Ronchetti,, 2009); see also Alfons et al., (2022) for the special case of mediation analysis. The non-robustness of ML estimation of the polychoric model has been demonstrated empirically by Foldnes & Grønneberg, (2020, 2022); Grønneberg & Foldnes, (2022) for the case of distributional misspecification. In this section, we introduce an estimator that is designed to be robust to partial misspecification when present, but remains (asymptotically) equivalent to the ML estimator of Olsson, (1979) when misspecification is absent. We furthermore derive the statistical properties of the proposed estimator.
Throughout this section, let $\{(X_{i},Y_{i})\}_{i=1}^{N}$ be an observed $N$-sized ordinal sample generated by discretizing latent variables $(\xi,\eta)$ that follow the unknown and unspecified mixture distribution $G_{\varepsilon}$ in (7). Hence, the polychoric model is possibly misspecified for an unknown fraction $\varepsilon$ of the sample.
@@ 5.1 The estimator
$f_{0}(x,y)=p_{xy}(\bm{\theta}_{*}).$
Conversely, if the polychoric model is misspecified $(\varepsilon>0)$, then $\widehat{f}_{N}(x,y)$ does not converge to the true $p_{xy}(\bm{\theta}_{*})$ because
$f_{\varepsilon}\left(x,y\right)\neq p_{xy}(\bm{\theta}_{*}),$
since we assume consequential misspecification.
It follows that if the polychoric model is misspecified, there exists no parameter value $\bm{\theta}\in\bm{\Theta}$ whose associated model probability $p_{xy}(\bm{\theta})$ the nonparametric estimate $\widehat{f}_{N}(x,y)$ converges to. Hence, it is indicative of model misspecification if $\widehat{f}_{N}(x,y)$ does not converge to any polychoric model probability $p_{xy}(\bm{\theta})$, resulting in a discrepancy between $\widehat{f}_{N}(x,y)$ and $p_{xy}(\bm{\theta})$. This observation can be exploited in model fitting by minimizing the discrepancy between the empirical relative frequencies, $\widehat{f}_{N}(x,y)$, and theoretical model probabilities, $p_{xy}(\bm{\theta})$, to find the most accurate fit that can be achieved with the polychoric model for the observed data. Specifically, our estimator minimizes with respect to $\bm{\theta}$ the loss function
@@
For the discrepancy function choice $\varphi(z)=\varphi^{\mathrm{MLE}}(z)=(z+1)\log(z+1)$, it can be easily verified that $\widehat{\bm{\theta}}_{N}$ coincides with the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ in (4). In the following, we motivate a specific choice of discrepancy function $\varphi(\cdot)$ that makes the estimator $\widehat{\bm{\theta}}_{N}$ less susceptible to misspecification of the polychoric model while preserving equivalence with ML estimation in the absence of misspecification.
@@
is referred to as Pearson residual (PR) (Lindsay,, 1994). It takes values in $[-1,+\infty)$ and can be interpreted as a goodness-of-fit measure. PR values close to 0 indicate a good fit between data and polychoric model at $\bm{\theta}$, whereas values toward $-1$ or +$\infty$+ ∞ indicate a poor fit because empirical response probabilities disagree with their model counterparts. To achieve robustness to misspecification of the polychoric model, responses that cannot be modeled well by the polychoric model, as indicated by their PR being away from 0, should receive less weight in the estimation procedure such that they do not over-proportionally affect the fit. Downweighting when necessary is achieved through a specific choice of discrepancy function $\varphi(\cdot)$ proposed by Welz, (2024), which is a special case of a function suggested by Ruckstuhl & Welsh, (2001). The discrepancy function reads
@@
where $c\in[0,\infty]$ is a prespecified tuning constant that governs the estimator’s behavior at the PR of each possible response. Figure 2 visualizes this function for the example choice $c=0.6$ as well as the ML discrepancy function function $\varphi^{\mathrm{MLE}}(z)=(z+1)\log(z+1)$. We stress that although the discrepancy function (11) can be negative, the loss function (9) is always nonnegative (Welz,, 2024).
@@
For the choice $c=+\infty$, minimizing the loss (9) is equivalent to maximizing the log-likelihood objective in (4), meaning that the estimator $\widehat{\bm{\theta}}_{N}$ is equal to $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ for this choice of $c$. More specifically, if a Pearson residual $z=\frac{\widehat{f}_{N}(x,y)}{p_{xy}(\bm{\theta})}-1$ of a response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ is such that $z\in[-1,c]$ for fixed $c\geq 0$, then the estimation procedure behaves at this response like in classic ML estimation. As argued before, in the absence of misspecification, $\widehat{f}_{N}(x,y)$ converges to $p_{xy}(\bm{\theta}_{*})$ for all responses $(x,y)\in\mathcal{X}\times\mathcal{Y}$, therefore all PRs are asymptotically equal to 0. Hence, if the polychoric model is correctly specified, then estimator $\widehat{\bm{\theta}}_{N}$ is asymptotically equivalent to the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ for any tuning constant value $c\geq 0$. On the other hand, if a response’s PR $z$ is larger than $c$, that is, $z>c\geq 0$, then the estimation procedure does not behave like in ML, but the response’s contribution to loss (9) is linear rather than super-linear like in ML (Figure 2). It follows that the influence of responses that cannot be fitted well by the polychoric model is downweighted to prevent them from dominating the fit. The tuning constant $c\geq 0$ is the “threshold” beyond which a PR will be downweighted, so the choice thereof determines what is considered an insufficient fit. The closer to 0 the tuning constant $c$ is chosen, the more robust the estimator.
@@ 5.2 Statistical properties
We first address what the generalized ML estimator $\widehat{\bm{\theta}}_{N}$ in (10) estimates. Its estimand is given by
@@
attains its global minimum of zero if and only if $\bm{\theta}=\bm{\theta}_{*}$, for any choice of $c\geq 0$. Thus, in the absence of misspecification, our estimator estimates the same quantity as the MLE, namely the true $\bm{\theta}_{*}$. We refer to Welz, (2024) for details.
Figure 3: The population estimand $\rho_{0}$ of the polychoric correlation coefficient for various degrees of misspecification fractions $\varepsilon$ ($x$-axis) and tuning constants $c$ (line colors). The ML estimand corresponds to $c=+\infty$. There are $K_{X}=K_{Y}=5$ response options and the true value corresponds to $\rho_{*}=0.5$ (dashed line). The misspecification type here is the same as in Figure 1.
However, in the presence of misspecification ($\varepsilon>0$), the estimand $\bm{\theta}_{0}$ is generally different from the true $\bm{\theta}_{*}$. The population estimand being different from the true value translates to biased estimates of the latter as a consequence of the misspecification. How much they are apart depends on the unknown fraction of misspecification $\varepsilon$, the unknown type of misspecification $H$, as well as the choice of tuning constant $c$ in $\varphi(\cdot)$. Mainly, the larger $\varepsilon$ (more severe misspecification) and $c$ (less downweighting of hard-to-fit responses), the further $\bm{\theta}_{0}$ is away from $\bm{\theta}_{*}$. Figure 3 illustrates this behavior for the polychoric correlation coefficient at an example misspecified distribution that is described in Section 4.1 and in which the true polychoric correlation under the correct model amounts to  $\rho_{*}=0.5$. For increasing misspecification fractions, the MLE ($c=+\infty)$ estimates a parameter value that is increasingly farther away from the true $\bm{\theta}_{*}$, where already a misspecification fraction of less than $\varepsilon=0.15$ suffices for a sign flip in the correlation coefficient. Conversely, choosing tuning constant $c$ to be near 0 results in a much less severe bias. For instance, even at misspecification degree $\varepsilon=0.3$, the difference between estimand and true value is less than approximately 0.1.
@@
A relevant question is whether the true parameter $\bm{\theta}_{*}$ can be recovered when $\varepsilon>0$ such that it can be estimated using $\widehat{\bm{\theta}}_{N}$ combined with a bias correction term. To derive such a bias correction term, one would need to impose assumptions on the misspecification fraction and type of misspecification. However, if one has strong prior beliefs about how the polychoric model is misspecified, modeling them explicitly rather than relying on the polychoric model seems more appropriate. Yet, one’s beliefs about misspecification may not be accurate, so attempts to explicitly model the misspecification may themselves result in a misspecified model. Consequently, robust estimation traditionally refrains from making assumptions on how a model may potentially be misspecified by leaving $\varepsilon$ and $H$ unspecified in mixture distribution (7). Instead, one may use a robust estimator to identify data points that cannot be modeled with the model at hand, like the one presented in this paper.
It can be shown that under certain regularity conditions that do not restrict the degree or type of possible partial misspecification beyond $\varepsilon\in[0,0.5)$, the robust estimator $\widehat{\bm{\theta}}_{N}$ is consistent for estimand $\bm{\theta}_{0}$ as well as asymptotically normally distributed. With this limit theory, one can construct standard errors and confidence intervals for every element in $\bm{\theta}_{0}$. Importantly, in the absence of model misspecification (such that $\bm{\theta}_{0}=\bm{\theta}_{*}$), the asymptotic covariance matrix of the robust estimator is equal to that of the fully efficient ML estimator as long as $c>0$. Hence, if the model is correctly specified, then robust estimator and ML estimator are asymptotically first and second order equivalent. We refer to Appendix A for a rigorous exposition and discussion of the robust estimator’s asymptotic properties.

### v4

@@ 4.3 Distributional misspecification
A model is distributionally misspecified when all observations in a given sample are generated by a distribution that is different from the model distribution. In the context of the polychoric model, this means that all ordinal observations are generated by a latent distribution that is nonnormal. Let $G$ denote the unknown nonnormal distribution that the latent variables $(\xi,\eta)$ jointly follow under distributional misspecification. The object of interest is the population correlation between latent $\xi$ and $\eta$ under distribution $G$, for which the normality-based MLE of Olsson, (1979) turns out to be substantially biased in many cases (e.g., Lyhagen & Ornstein,, 2023; Foldnes & Grønneberg,, 2020, 2022; Jin & Yang-Wallentin,, 2017). As such, distributional misspecification is fundamentally different from partial misspecification: In the former, one attempts to estimate the population correlation of the nonnormal and unknown distribution that generated a sample, instead of estimating the polychoric correlation coefficient (which is the correlation under standard bivariate normality). In the latter, one attempts to estimate the polychoric correlation coefficient with a contaminated sample that has only been partly generated by latent normality (that is, the polychoric model). The assumption that the polychoric model is only partially misspecified for some uninformative observations enables one to still estimate the polychoric correlation coefficient of that model, which would not be feasible under distributional misspecification (at least not without additional assumptions).
Despite not being designed for distributional misspecification, the robust estimator introduced in the next section can offer a robustness gain in some situations where the polychoric model is distributionally misspecified. We discuss this in more detail in Section 8.
@@ 5 Robust estimation of polychoric correlation
The behavior of ML estimates of any model crucially depends on correct specification of that model. Indeed, ML estimation can be severely biased even when the assumed model is only slightly misspecified (e.g., Huber & Ronchetti,, 2009; Hampel et al.,, 1986; Huber,, 1964). For instance, in many models of continuous variables like regression models, one single observation from a different distribution can be enough to make the ML estimator converge to an arbitrary value (Huber & Ronchetti,, 2009; see also Alfons et al.,, 2022, for the special case of mediation analysis). The non-robustness of ML estimation of the polychoric model has been demonstrated empirically by Foldnes & Grønneberg, (2020, 2022); Grønneberg & Foldnes, (2022) for the case of distributional misspecification. In this section, we introduce an estimator that is designed to be robust to partial misspecification when present, but remains (asymptotically) equivalent to the ML estimator of Olsson, (1979) when misspecification is absent. We furthermore derive the statistical properties of the proposed estimator.
Throughout this section, let $\{(X_{i},Y_{i})\}_{i=1}^{N}$ be an observed ordinal sample of size $N$ generated by discretizing latent variables $(\xi,\eta)$ that follow the unknown and unspecified contaminated distribution $G_{\varepsilon}$ in (7). Hence, the polychoric model is possibly misspecified for an unknown fraction $\varepsilon$ of the sample.
@@ 5.1 The estimator
$f_{0}(x,y)=p_{xy}(\bm{\theta}_{*}),$
for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Conversely, if the polychoric model is misspecified $(\varepsilon>0)$, then $\widehat{f}_{N}(x,y)$ may not converge to the true $p_{xy}(\bm{\theta}_{*})$ because
$f_{\varepsilon}\left(x,y\right)\neq p_{xy}(\bm{\theta}_{*})$
for some $(x,y)\in\mathcal{X}\times\mathcal{Y}$, since we assume consequential misspecification.
It follows that if the polychoric model is misspecified, there exists no parameter value $\bm{\theta}\in\bm{\Theta}$ for which the nonparametric estimates $\widehat{f}_{N}(x,y)$ converge pointwise to the associated model probabilities $p_{xy}(\bm{\theta})$ for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Hence, it is indicative of model misspecification if there exists at least one response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ for which $\widehat{f}_{N}(x,y)$ does not converge to any polychoric model probability $p_{xy}(\bm{\theta})$, resulting in a discrepancy between $\widehat{f}_{N}(x,y)$ and $p_{xy}(\bm{\theta})$. [^7] Actually, if said convergence fails, it does so for at least two responses. Due to the relative frequencies summing up to 1, if the convergence fails for one response, it must also fail for at least one more. This observation can be exploited in model fitting by minimizing the discrepancy between the empirical relative frequencies, $\widehat{f}_{N}(x,y)$, and theoretical model probabilities, $p_{xy}(\bm{\theta})$, to find the most accurate fit that can be achieved with the polychoric model for the observed data. Specifically, our estimator minimizes with respect to $\bm{\theta}$ the loss function
@@
For the choice of discrepancy function $\varphi(z)=\varphi^{\mathrm{MLE}}(z)=(z+1)\log(z+1)$, it can be easily verified that $\widehat{\bm{\theta}}_{N}$ coincides with the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ in (4). In the following, we motivate a specific choice of discrepancy function $\varphi(\cdot)$ that makes the estimator $\widehat{\bm{\theta}}_{N}$ less susceptible to misspecification of the polychoric model while preserving equivalence with ML estimation in the absence of misspecification.
@@
is referred to as Pearson residual (PR) (Lindsay,, 1994). It takes values in [1,+)[-1,$\infty$)[ - 1 , + ∞ ) and can be interpreted as a goodness-of-fit measure. PR values close to 0 indicate a good fit between data and polychoric model at $\bm{\theta}$, whereas values toward 1-1- 1 or +$\infty$+ ∞ indicate a poor fit because empirical response probabilities disagree with their model counterparts. To achieve robustness to misspecification of the polychoric model, responses that cannot be modeled well by the polychoric model, as indicated by their PR being away from 0, should receive less weight in the estimation procedure such that they do not over-proportionally affect the fit. Downweighting when necessary is achieved through a specific choice of discrepancy function $\varphi(\cdot)$ proposed by Welz, (2024), which is a special case of a function suggested by Ruckstuhl & Welsh, (2001). The discrepancy function reads
@@
where $c\in[0,\infty]$ is a prespecified tuning constant that governs the estimator’s behavior at the PR of each possible response. Figure 2 visualizes this function for the example choice $c=0.6$ as well as the ML discrepancy function $\varphi^{\mathrm{MLE}}(z)=(z+1)\log(z+1)$. Note that deducting 1 in (9) and adding it again in (11) is purely for keeping the interpretation that a PR close to 0 indicates a good fit. We further stress that although the discrepancy function (11) can be negative, the loss function (9) is always nonnegative (Welz,, 2024).
@@
For the choice $c=+\infty$, minimizing the loss (9) is equivalent to maximizing the log-likelihood objective in (4), meaning that the estimator $\widehat{\bm{\theta}}_{N}$ is equal to $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ for this choice of $c$. More specifically, if a Pearson residual $z=\frac{\widehat{f}_{N}(x,y)}{p_{xy}(\bm{\theta})}-1$ of a response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ is such that $z\in[-1,c]$ for fixed $c\geq 0$, then the estimation procedure behaves at this response like in classic ML estimation. As argued before, in the absence of misspecification, $\widehat{f}_{N}(x,y)$ converges to $p_{xy}(\bm{\theta}_{*})$ for all responses $(x,y)\in\mathcal{X}\times\mathcal{Y}$, therefore all PRs are asymptotically equal to 0. Hence, if the polychoric model is correctly specified, then estimator $\widehat{\bm{\theta}}_{N}$ is asymptotically equivalent to the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ for any tuning constant value $c\geq 0$. On the other hand, if a response’s PR $z$ is larger than $c$, that is, $z>c\geq 0$, then the estimation procedure does not behave like in ML, but the response’s contribution to loss (9) is linear rather than super-linear like in ML (Figure 2). It follows that the influence of responses that cannot be fitted well by the polychoric model is downweighted to prevent them from dominating the fit. The tuning constant $c\geq 0$ is the threshold beyond which a PR will be downweighted, so the choice thereof determines what is considered an insufficient fit. The closer to 0 the tuning constant $c$ is chosen, the more robust the estimator is in theory. In Appendix C, we explore different values of $c$ in simulations and motivate a specific choice that we use for all numerical results in this paper, namely $c=0.6$.
[…]
@@ 5.2 Statistical properties
We first address what quantity is estimated by the proposed estimator before discussing its asymptotic behavior.
5.2.1 Estimand
The estimand of the estimator $\widehat{\bm{\theta}}_{N}$ in (10) is given by
@@ 5.2.1 Estimand
attains its global minimum of zero if and only if $\bm{\theta}=\bm{\theta}_{*}$, for any choice of $c\geq 0$. Thus, in the absence of contamination, our estimator estimates the same quantity as the MLE, namely the true $\bm{\theta}_{*}$. In other words, it obtains the true $\bm{\theta}_{*}$ in population when the model is correctly specified, a property known as Fisher consistency. We refer to Welz, (2024) for details.
Figure 3: The population estimand $\rho_{0}$ of the polychoric correlation coefficient for various degrees of contamination fractions $\varepsilon$ ($x$-axis) and tuning constants $c$ (line colors), for the same contamination distribution as in Figure 1. The ML estimand corresponds to $c=+\infty$. There are $K_{X}=K_{Y}=5$ response options and the true value corresponds to $\rho_{*}=0.5$ (dashed line).
However, in the presence of misspecification ($\varepsilon>0$), the sampling distribution differs from the model distribution such that the estimand $\bm{\theta}_{0}$—the parameter that minimizes the loss evaluated at the sampling distribution—is generally different from the true $\bm{\theta}_{*}$ (cf. White,, 1982). [^8] White, (1982) is concerned with (quasi-)maximum likelihood estimation, but the same reasoning applies to our estimator. In the context of White, (1982), the ML estimand of a misspecified model corresponds to the parameter value minimizing the Kullback-Leibler divergence between the sampling density and the model density. For $c=\infty$ (the ML case), the population analogue of our loss function (9) corresponds to the Kullback-Leibler divergence between the sampling density $f_{\varepsilon}\left(x,y\right)$ and the model density $p_{xy}(\cdot)$, that is, $\bm{\theta}\mapsto L\left(\bm{\theta},f_{\varepsilon}\right)$. For finite choices of $c$, our loss function is designed to yield an estimand that is closer to the true $\bm{\theta}_{*}$ than that of the ML loss by downweighting responses that cannot be sufficiently well modeled. The population estimand being different from the true value translates to biased estimates of the latter as a consequence of the misspecification.
How much the estimand $\bm{\theta}_{0}$ differs from the true $\bm{\theta}_{*}$ depends on the unknown fraction of contamination $\varepsilon$, the unknown type of contamination $H$, as well as the choice of tuning constant $c$ in $\varphi(\cdot)$. Mainly, the larger $\varepsilon$ (more severe misspecification) and $c$ (less downweighting of hard-to-fit responses), the further $\bm{\theta}_{0}$ is away from $\bm{\theta}_{*}$. Figure 3 illustrates this behavior for the polychoric correlation coefficient at an example misspecified distribution that is described in Section 4.2 and in which the true polychoric correlation under the correct model amounts to  $\rho_{*}=0.5$. For increasing contamination fractions, the MLE ($c=+\infty)$ estimates a parameter value that is increasingly farther away from the true $\bm{\theta}_{*}$, where already a contamination fraction of less than $\varepsilon=0.15$ suffices for a sign flip in the correlation coefficient. Conversely, choosing tuning constant $c$ to be near 0 results in a much less severe bias. For instance, even at contamination fraction $\varepsilon=0.4$, the difference between estimand and true value is approximately 0.1 or less.
@@
A relevant question is whether the true parameter $\bm{\theta}_{*}$ can be recovered when $\varepsilon>0$ such that it can be estimated using $\widehat{\bm{\theta}}_{N}$ combined with a bias correction term. To derive such a bias correction term, one would need to impose assumptions on the contamination fraction and type of contamination. However, if one has strong prior beliefs about how the polychoric model is misspecified, modeling them explicitly rather than relying on the polychoric model seems more appropriate. Yet, one’s beliefs about misspecification may not be accurate, so attempts to explicitly model the misspecification may themselves result in a misspecified model. Consequently, robust estimation traditionally refrains from making assumptions on how a model may potentially be misspecified by leaving $\varepsilon$ and $H$ unspecified in the contaminated distribution (7). Instead, one may use a robust estimator to identify data points that cannot be modeled with the model at hand, like the one presented in this paper.
5.2.2 Asymptotic analysis
It can be shown that under certain standard regularity conditions that do not restrict the degree or type of possible partial misspecification beyond $\varepsilon\in[0,0.5)$, the robust estimator $\widehat{\bm{\theta}}_{N}$ is consistent for estimand $\bm{\theta}_{0}$ as well as asymptotically normally distributed.Specifically, under said regularity conditions and fixed tuning constant $c>0$, Theorem 1 in Appendix A establishes that
as well as
as $N\to\infty$, where “$\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}$” and “$\stackrel{{\scriptstyle\mathrm{d}}}{{\longrightarrow}}$” denote convergence in probability and distribution, respectively. The asymptotic covariance matrix has a sandwich-type construction
$\bm{\Sigma}\left({\bm{\theta}}\right)=\bm{M}\left(\bm{\theta}\right)^{-1}\bm{U}\left(\bm{\theta}\right)\bm{M}\left(\bm{\theta}\right)^{-1},\qquad\bm{\theta}\in\bm{\Theta},$
where the $d\times d$ matrices
$\bm{M}\left(\bm{\theta}\right)=\frac{\partial^{2}L\big{(}\bm{\theta},f_{\varepsilon}\big{)}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}\quad\ \textnormal{and}\ \quad\bm{U}\left(\bm{\theta}\right)=\mathbb{V}\mathrm{ar}_{f_{\varepsilon}}\left[\frac{\partial\log\left(p_{XY}(\bm{\theta})\right)}{\partial\bm{\theta}}\mathds{1}\left\{\frac{f_{\varepsilon}\left(X,Y\right)}{p_{XY}}-1\in[-1,c]\right\}\right],$
respectively, are the Hessian matrix of the population loss and the covariance matrix (evaluated at $f_{\varepsilon}$) of the likelihood score function—that is, the gradient of $\log(p_{XY}(\bm{\theta}))$—weighted by stochastic binary weights whether the Pearson residual is smaller than or equal to the tuning constant $c$ [^9] It is worth mentioning that for the MLE ($c=+\infty$), our sandwich-type covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ reduces to that of White, (1982) and Huber, (1967), who studied the limit distribution of ML in misspecified models. We derive closed-form expressions of the matrices $\bm{M}\left(\bm{\theta}\right)$ and $\bm{U}\left(\bm{\theta}\right)$ as well as their properties in Appendix A.
The asymptotic covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ of our estimator is unobserved in practice because it depends on the unknown quantities $\bm{\theta}_{0}$ and $f_{\varepsilon}$. Yet, $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ can be consistently estimated by replacing $\bm{\theta}_{0}$ and $f_{\varepsilon}$ by their corresponding consistent estimators $\widehat{\bm{\theta}}_{N}$ and $\widehat{f}_{N}$, respectively. Details are provided in Appendix A.
With this limit theory, one can construct standard errors and confidence intervals for every element in $\bm{\theta}_{0}$. Importantly, in the absence of contamination (such that $\bm{\theta}_{0}=\bm{\theta}_{*}$), the asymptotic covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ of the robust estimator reduces to that of the fully efficient MLE (i.e., inverse Fisher information matrix) as long as $c>0$. It follows that there is no loss of statistical efficiency when there is no contamination. [^10] This property does not contradict standard theory on the Cramér-Rao lower bound (CRLB) for the variance of unbiased estimators (e.g., Theorem 7.3.9 in Casella & Berger,, 2002). In brief, if a statistical model is correctly specified, the variance of the MLE asymptotically attains this lower bound such that no unbiased estimator can be asymptotically more efficient than the MLE. Nevertheless, this does not imply that no other estimators can (asymptotically) attain the CRLB. When the polychoric model is correctly specified (i.e., no contamination), our estimator has the same asymptotic variance as the MLE, which satisfies the Cramér-Rao lower bound with equality. Hence, if the polychoric model is correctly specified, the robust estimator and the MLE are asymptotically first and second order equivalent. We refer to Appendix A for a rigorous exposition and discussion of the robust estimator’s asymptotic properties.